One side of triangle is 2 times the shortest side. The third side is 16 feet more than the shortest side. The perimeter is 80. Find all three sides.
Answer by aliafzal14(13)

ACCORDING TO THE GIVEN CONDITION
SHORTEST SIDE IS EQUAL TO THE = A= X
2ND SIDE IS EQUAL TO = B =2X
3RD SIDE IS EQUAL TO = C =16 + X
IN MATHEMATICALLY,
PERIMETER = A +B +C
PERIMETER = X + +2X + (16 +X)
80 = 4X + 16
4X = 80-16
4X = 64
X = 64/4
X = 16
SHORTEST SIDE IS EQUAL TO THE = A= X = 16
2ND SIDE IS EQUAL TO = B =2X = 2*16 = 32
3RD SIDE IS EQUAL TO = C =16 + X = 16 + 16 = 32